# Instructions for Authors: Title in Title Case

### v1

@@ B.2.2 Baseline Settings
B-IRL: Following [Hüyük et al., 2022], we have run Metropolis-Hastings algorithm for 10,000 iterations to obtain 1,000 samples on synthetic dataset (500 samples and 5,000 iterations for real-world ML-100K dataset) with intervals of 10 iterations between each sample after 10,000 burn-in iterations (5,000 for real-world ML-100K dataset). The final estimated parameters are formed by simply averaging all 1,000 samples. (500 for real-world ML-100K dataset).
B-ICB: We followed the setting used in [Hüyük et al., 2022]’s work. We changed the setting of $\mathrm{compute\_rhox}$ period (i.e., computing the reward parameters) in real-world ML-100K dataset experiments. Since that ML-100K’s user embedding used one-hot encoding, which may make the final matrix $\bm{\beta}_{n}$ to concatenate in $\mathrm{compute\_rhox}$ be non-invertible, we replace the original $\mathrm{np.linalg.inv}$ method with $\mathrm{np.linalg.pinv}$ method to compute the pseudo inverse of $\bm{\beta}_{n}$, making sure experiment can be done without computational errors. And due to that $\bm{\beta}_{n}$ may be non-invertible, we also dismiss the $\mathrm{det}(\beta_{n}/\sigma^{2})$ during iteration, where $\mathrm{det}(\bm{A})$ means the determinant of matrix $\bm{A}$.
@@ B.2.3 IBCB Settings
B.3 Details of the Evaluation Protocol
OL-Fitness: OL-data’s candidate context sets, sequences and reward parameters keep same during the comparison by using fixed seeds. Note that BC algorithms cannot obtain reward parameters, and thus were not included in this comparison scope. But in dup-test period, we used ‘train log behaviors fitness’ to evaluate whether BC algorithm suffers from contradictory data.
BT-Fitness: BT-data’s candidate context sets, sequences and reward parameters keep same during the comparison by using fixed seeds, and expert do not update its policy parameters in BT-data phase.
B.4 Results Details
@@
All results were averaged over 5 different runs, and sampled the data over 5 different random seeds for each train data log (behavioral evolution history log) of expert on both synthetic and real-world ML-100K dataset.
B.4.1 Results of Out-of-distribution Data on Synthetic Dataset
@@
Table 10: Expert policy (SBUCB) parameters’ estimation (Out-of-distribution data). Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness (OOD)
std=0 std=0.03 std=0.07 std=0.1
@@
B-IRL 0.802±0.005 0.793±0.013 0.795±0.015 0.793±0.016
BaseSVC 0.764±0.009 0.756±0.021 0.756±0.022 0.750±0.023
B-ICB-S 0.775±0.039 0.738±0.065 0.755±0.050 0.754±0.025
IBCB (Ours) 0.834±0.012 0.822±0.031 0.821±0.033 0.816±0.029
Algorithm Batch Test Average Reward (OOD)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 0.709±0.015 0.695±0.036 0.694±0.031 0.684±0.030
B-IRL 0.642±0.019 0.622±0.047 0.622±0.042 0.608±0.041
BaseSVC 0.627±0.021 0.607±0.052 0.606±0.046 0.591±0.046
B-ICB-S 0.632±0.008 0.600±0.030 0.607±0.032 0.596±0.032
IBCB (Ours) 0.654±0.021 0.632±0.054 0.632±0.048 0.618±0.046
Table 11: Expert policy (SBTS) parameters’ estimation (Out-of-distribution data). Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness (OOD)
std=0 std=0.03 std=0.07 std=0.1
SBTS 1 1 1 1
B-IRL 0.792±0.005 0.795±0.010 0.793±0.006 0.793±0.005
BaseSVC 0.767±0.015 0.782±0.017 0.780±0.019 0.776±0.014
B-ICB-S 0.714±0.065 0.710±0.096 0.712±0.081 0.724±0.086
IBCB (Ours) 0.839±0.022 0.859±0.021 0.858±0.022 0.852±0.022
Algorithm Batch Test Average Reward (OOD)
std=0 std=0.03 std=0.07 std=0.1
SBTS 0.719±0.028 0.739±0.032 0.735±0.032 0.729±0.032
B-IRL 0.650±0.034 0.676±0.037 0.671±0.039 0.663±0.037
BaseSVC 0.640±0.041 0.670±0.045 0.666±0.047 0.658±0.044
B-ICB-S 0.620±0.028 0.643±0.028 0.642±0.026 0.637±0.027
IBCB (Ours) 0.667±0.040 0.697±0.043 0.693±0.044 0.684±0.044
As shown in Table 10 and Table 11, IBCB still outperforms other IL baselines in expert policy’s parameter estimation on Out-of-distribution data.
B.4.2 Detailed Results of Contradictory Data Testing on Synthetic Dataset
Table 12: Contradictory (duplicated) data (appeared at expert’s OL-data behavioral evolution history logs) on synthetic dataset testing. $\mathrm{dup}$ denotes the quantity of duplicated episodes in OL-data phase.
algorithm Online Train Log Fitness ($\mathrm{dup}<5$)
dup=0 dup=1 dup=2 dup=3 dup=4
SBUCB 0.879±0.016 0.880±0.015 0.880±0.015 0.879±0.015 0.880±0.018
B-IRL 0.808±0.039 0.808±0.037 0.808±0.038 0.806±0.038 0.807±0.038
IBCB (Ours) 0.870±0.028 0.870±0.03 0.868±0.029 0.867±0.030 0.869±0.030
Algorithm Train Log Behaviors Fitness (BC) ($\mathrm{dup}<5$)
dup=0 dup=1 dup=2 dup=3 dup=4
BaseSVC
0.823±0.026 0.825±0.025 0.824±0.026 0.821±0.026 0.820±0.026
Algorithm Online Train Log Fitness ($\mathrm{dup}\geq 5$)
dup=5 dup=6 dup=7 dup=8 dup=9
SBUCB 0.879±0.015 0.880±0.016 0.879±0.015 0.879±0.015 0.878±0.017
B-IRL 0.807±0.037 0.809±0.037 0.810±0.037 0.809±0.037 0.808±0.037
IBCB (Ours) 0.868±0.029 0.870±0.030 0.872±0.029 0.870±0.028 0.869±0.029
Algorithm Train Log Behaviors Fitness (BC) ($\mathrm{dup}\geq 5$)
dup=5 dup=6 dup=7 dup=8 dup=9
BaseSVC
0.819±0.025 0.819±0.025 0.819±0.024 0.817±0.023 0.816±0.023
Note that BaseSVC cannot obtain reward parameters, and thus were not included in OL-Fitness. We used ‘Train Log Behaviors Fitness’ to evaluate whether BaseSVC suffers from contradictory data and make comparison with IBCB and other baselines. As shown in Table 12, contradictory data (same context in different time periods) undermines BaseSVC’s train log behaviors fitness, but IBCB stayed stable in OL-Fitness.
B.4.3 Results of IBCB’s Parameter Test on Synthetic Dataset
Table 13: Reward parameters’ estimation and training on synthetic dataset (IBCB’s parameter test). Experts do not need train since they produce train logs (behavioral evolution history logs) for baselines and IBCB. BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison. Original Expert is SBUCB.
Algorithm Online Train Log Fitness (Param Test)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 0.883±0.007 0.879±0.016 0.878±0.018 0.874±0.019
B-IRL 0.827±0.012 0.808±0.039 0.806±0.037 0.802±0.036
BaseSVC / / / /
B-ICB-S 0.805±0.020 0.771±0.031 0.780±0.034 0.771±0.024
IBCB-$\alpha$-1 0.882±0.012 0.870±0.028 0.871±0.033 0.865±0.029
IBCB-$\alpha$-0.6 0.874±0.011 0.863±0.032 0.861±0.038 0.854±0.033
IBCB-$\alpha$-0.8 0.880±0.012 0.868±0.030 0.868±0.034 0.862±0.030
IBCB-$\alpha$-1.2 0.884±0.012 0.872±0.028 0.873±0.032 0.866±0.029
IBCB-$\alpha$-1.5 0.886±0.011 0.872±0.029 0.873±0.033 0.867±0.028
IBCB-$\alpha$-2 0.886±0.012 0.874±0.026 0.875±0.032 0.869±0.028
Algorithm Train Time(sec.) (Param Test)
std=0 std=0.03 std=0.07 std=0.1
SBUCB / / / /
B-IRL 36.997±0.803 37.670±1.328 39.910±3.419 37.424±0.563
BaseSVC 116.86±20.63 116.39±26.58 132.18±19.86 117.25±19.47
B-ICB-S 86.365±0.827 87.599±4.489 86.85±1.162 85.713±0.676
IBCB-$\alpha$-1 6.622±2.621 7.992±3.218 7.186±1.577 8.615±4.599
IBCB-$\alpha$-0.6 6.867±2.321 6.802±2.731 7.140±2.044 7.495±4.268
IBCB-$\alpha$-0.8 6.912±2.377 7.913±3.366 7.118±1.762 8.783±5.314
IBCB-$\alpha$-1.2 7.252±2.970 9.823±3.789 7.745±1.758 10.90±6.446
IBCB-$\alpha$-1.5 1.494±0.380 1.453±0.442 1.412±0.359 1.858±0.679
IBCB-$\alpha$-2 1.806±0.465 1.423±0.494 1.620±0.480 1.864±0.686
Table 14: Expert policy parameters’ estimation on synthetic dataset (IBCB’s parameter test). Expert policy produce same actions compared with itself, so BT-Fitness of expert policy is 1. Original Expert is SBUCB.
Algorithm Batch Test Log Fitness (Param Test)
std=0 std=0.03 std=0.07 std=0.1
SBUCB 1 1 1 1
B-IRL 0.828±0.005 0.820±0.013 0.820±0.014 0.818±0.016
BaseSVC 0.791±0.008 0.781±0.020 0.781±0.020 0.777±0.002
B-ICB-S 0.800±0.032 0.767±0.055 0.780±0.045 0.776±0.022
IBCB-$\alpha$-1 0.856±0.011 0.843±0.028 0.842±0.030 0.838±0.027
IBCB-$\alpha$-0.6 0.856±0.011 0.843±0.029 0.842±0.030 0.838±0.027
IBCB-$\alpha$-0.8 0.856±0.011 0.843±0.028 0.842±0.030 0.838±0.027
IBCB-$\alpha$-1.2 0.856±0.011 0.843±0.028 0.842±0.030 0.838±0.027
IBCB-$\alpha$-1.5 0.859±0.011 0.842±0.028 0.843±0.031 0.838±0.027
IBCB-$\alpha$-2 0.858±0.010 0.843±0.027 0.843±0.032 0.839±0.027
Algorithm Batch Test Average Reward (Param Test)
std=0 std=0.03 std=0.07 std=0.1

### v2

@@ 5.4.2 Baseline Settings
B-IRL: Following [15], we have run Metropolis-Hastings algorithm for 10,000 iterations to obtain 1,000 samples on synthetic dataset (500 samples and 5,000 iterations for real-world ML-100K dataset) with intervals of 10 iterations between each sample after 10,000 burn-in iterations (5,000 for real-world ML-100K dataset). The final estimated parameters are formed by simply averaging all 1,000 samples. (500 for real-world ML-100K dataset).
B-ICB: We followed the setting used in [15]’s work. We changed the setting of $\mathrm{compute\_rhox}$ period (i.e., computing the reward parameters) in real-world ML-100K dataset experiments. Since that ML-100K’s user embedding used one-hot encoding, which may make the final matrix $\bm{\beta}_{n}$ to concatenate in $\mathrm{compute\_rhox}$ be non-invertible, we replace the original $\mathrm{np.linalg.inv}$ method with $\mathrm{np.linalg.pinv}$ method to compute the pseudo inverse of $\bm{\beta}_{n}$, making sure experiment can be done without computational errors. And due to that $\bm{\beta}_{n}$ may be non-invertible, we also dismiss the $\mathrm{det}(\beta_{n}/\sigma^{2})$ during iteration, where $\mathrm{det}(\bm{A})$ means the determinant of matrix $\bm{A}$.
@@ 5.4.3 IBCB Settings
5.5 Experimental Results & Analyses
In this section, we compared several indicators between IBCB with other baselines. We used data sampling method to train B-ICB (namely B-ICB-S, where ‘-S’ means ‘used sampled data’) since B-ICB has not yet been adapted to batched version, so B-ICB can only be trained in environment that quantity of train data is not so large.
@@ 5.5 Experimental Results & Analyses
All results were averaged over 5 different runs, and sampled the data over 5 different random seeds for each train data log (behavioral evolution history log) of expert on both synthetic and real-world ML-100K dataset. The source code has been available at https://anonymous.4open.science/r/IBCB-F25F.
5.5.1 Experiments on Synthetic Dataset
TABLE I: Reward parameters’ estimation (Upper Subtable) and Train time (Lower Subtable) comparison on synthetic dataset, where $\mathrm{std}$ denotes the standard deviation of zero-mean Gaussian noise added to the final reward. Experts (SBUCB & SBTS) do not need to train since they produce the train logs (behavioral evolution history logs) for baselines and IBCB.
Algorithm Online Train Log Fitness (OL-Fitness) Algorithm Online Train Log Fitness (OL-Fitness)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 0.883±0.007 0.879±0.016 0.878±0.018 0.874±0.019 SBTS 0.857±0.017 0.876±0.012 0.872±0.014 0.871±0.008
B-IRL 0.827±0.012 0.808±0.039 0.806±0.037 0.802±0.036 B-IRL 0.457±0.020 0.468±0.029 0.480±0.023 0.476±0.039
B-ICB-S 0.805±0.020 0.771±0.031 0.780±0.034 0.771±0.024 B-ICB-S 0.424±0.020 0.451±0.036 0.457±0.016 0.454±0.024
IBCB (Ours) 0.868±0.008 0.858±0.036 0.859±0.038 0.861±0.029 IBCB (Ours) 0.835±0.014 0.856±0.009 0.849±0.014 0.849±0.009
Algorithm Train Time (sec.) Algorithm Train Time (sec.)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB / / / / SBTS / / / /
B-IRL 36.997±0.803 37.67±1.328 39.91±3.419 37.424±0.563 B-IRL 38.46±3.550 36.419±0.187 37.843±3.486 43.469±6.228
BaseSVC 116.86±20.63 116.39±26.58 132.18±19.86 117.25±19.47 BaseSVC 96.603±9.619 87.310±8.435 89.240±6.566 90.472±8.494
B-ICB-S 86.365±0.827 87.599±4.489 86.85±1.162 85.713±0.676 B-ICB-S 91.166±1.164 89.193±1.477 89.892±1.162 89.483±1.585
IBCB (Ours) 0.815±0.045 0.999±0.128 0.866±0.137 1.016±0.109 IBCB (Ours) 0.778±0.035 0.809±0.077 0.772±0.034 0.871±0.216
@@ 5.5.1 Experiments on Synthetic Dataset
TABLE II: Expert policy (SBUCB) parameters’ estimation (Upper Subtable), Expert policy (SBTS) parameters’ estimation (Lower Subtable) on synthetic dataset. Expert policy produces same actions compared with itself, so Batch Test Log Fitness (BT-Fitness) of expert policy is 1.
Algorithm Batch Test Log Fitness (BT-Fitness) Algorithm Batch Test Average Reward (BT-AR)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 1 1 1 1 SBUCB 0.641±0.010 0.629±0.028 0.630±0.024 0.622±0.024
B-IRL 0.828±0.005 0.820±0.013 0.820±0.014 0.818±0.016 B-IRL 0.588±0.015 0.571±0.037 0.572±0.034 0.561±0.034
BaseSVC 0.791±0.008 0.781±0.020 0.781±0.020 0.777±0.020 BaseSVC 0.577±0.017 0.559±0.042 0.559±0.037 0.547±0.037
B-ICB-S 0.800±0.032 0.767±0.055 0.780±0.045 0.776±0.022 B-ICB-S 0.581±0.006 0.553±0.024 0.560±0.025 0.552±0.025
IBCB (Ours) 0.972±0.011 0.963±0.031 0.964±0.028 0.972±0.016 IBCB (Ours) 0.647±0.012 0.638±0.026 0.639±0.019 0.628±0.022
Algorithm Batch Test Log Fitness (BT-Fitness) Algorithm Batch Test Average Reward (BT-AR)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBTS 1 1 1 1 SBTS 0.650±0.020 0.662±0.025 0.661±0.025 0.656±0.024
B-IRL 0.816±0.007 0.820±0.015 0.816±0.008 0.816±0.006 B-IRL 0.595±0.026 0.615±0.028 0.611±0.031 0.605±0.030
BaseSVC 0.789±0.011 0.807±0.009 0.801±0.015 0.798±0.012 BaseSVC 0.586±0.032 0.611±0.035 0.607±0.037 0.600±0.034
B-ICB-S 0.740±0.059 0.742±0.089 0.741±0.074 0.752±0.076 B-ICB-S 0.571±0.021 0.589±0.022 0.588±0.021 0.584±0.023
IBCB (Ours) 0.963±0.012 0.958±0.010 0.958±0.006 0.958±0.005 IBCB (Ours) 0.651±0.027 0.667±0.027 0.666±0.026 0.658±0.029
Table I display OL-Fitness and training time for experts (SBUCB & SBTS), baselines, and IBCB. IBCB outperforms others with higher OL-Fitness and faster training, thanks to its optimal reward parameter estimation and OSQP solver’s efficiency in managing large-scale QP problems with nearly linear complexity. BaseSVC trains slower due to its complexity near to quadratic, while B-IRL requires many Metropolis-Hastings iterations for better performance. B-ICB’s quadratic complexity slows its training, despite sampling data.
Table II show that IBCB also achieves higher BT-Fitness and BT-AR, highlighting the effectiveness of its reward imputation method for learning from the expert’s behavioral evolution history. This confirms IBCB’s ability to learn the expert policy’s parameters effectively and efficiently, even in out-of-distribution (OOD) scenarios as seen in Table VI in Section 5.6.1, demonstrating its robustness to distribution shifts.
5.5.2 Experiments on Real-World Dataset (ML-100K Dataset)
TABLE III: Reward parameters’ estimation , Train time comparison and Expert policy parameters’ estimation on ML-100K dataset. Experts do not need to train since they produce train logs (behavioral evolution history logs) for baselines and IBCB. BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison. Expert policy produces same actions compared with itself, so BT-Fitness of expert policy is 1.
Algorithm OL-Fitness Train Time (sec.) BT-Fitness BT-AR Algorithm OL-Fitness Train Time (sec.) BT-Fitness BT-AR
SBUCB 0.891±0.006 / 1 0.193±0.005 SBTS 0.868±0.015 / 1 0.200±0.005
B-IRL 0.830±0.010 43.669±1.282 0.864±0.008 0.189±0.005 B-IRL 0.851±0.015 47.005±4.311 0.870±0.030 0.199±0.007
BaseSVC / 162.767±12.135 0.804±0.012 0.189±0.005 BaseSVC / 192.83±33.199 0.801±0.019 0.198±0.007
B-ICB-S 0.746±0.032 147.474±4.633 0.767±0.028 0.187±0.007 B-ICB-S 0.721±0.030 145.441±5.916 0.759±0.035 0.196±0.004
IBCB (Ours) 0.862±0.023 4.036±0.111 0.925±0.019 0.192±0.005 IBCB (Ours) 0.869±0.021 8.830±1.825 0.956±0.006 0.201±0.005
Results on ML-100K dataset of different indicators are reported in Table III. Similar to the results on synthetic dataset, IBCB achieved significantly higher performance than other baselines in all indicators. From the outstanding results of training speed and average reward, we concluded that IBCB can also adapt to learning real-world behavioral evolution history generated by expert with great efficiency.
5.5.3 Experiments on Fairness-Aware Expert Limitation
In the original fairness-aware experiment, we set the top-$k$ with $k=2$, which means we choose the final action from the top-$2$ based on normalized probabilities.
TABLE IV: Fairness comparison (Upper Subtable), Fairness-aware train time comparison (Lower Subtable) on synthetic dataset (CFR definition can be found in Section 5.3.4). Expert selects action from top-$\bm{2}$ actions with proportion of its learned probability distribution through all actions. Expert is defined as the fairest policy from the CFR metric, so its CFR is always 0. Fairness-aware experts (SBUCB & SBTS) do not need to train since they produce the train logs (behavioral evolution history logs) for baselines and IBCB.
Algorithm Cumulative Fairness Regret (CFR) Algorithm Cumulative Fairness Regret (CFR)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 0 0 0 0 SBTS 0 0 0 0
B-IRL 2175.0±159.1 2330.0±235.8 2347.2±305.4 2449.0±344.1 B-IRL 2106.5±136.6 2080.7±236.3 2102.9±227.4 2136.9±192.1
BaseSVC 2027.5±184.2 2218.2±342.6 2248.0±402.2 2367.8±449.9 BaseSVC 1848.8±232.4 1776.9±427.0 1763.8±391.5 1841.2±344.7
B-ICB-S 2162.3±205.5 2661.7±169.6 2507.1±240.8 2437.6±175.9 B-ICB-S 2679.5±405.6 2922.5±451.3 2660.4±305.1 2800.2±197.7
IBCB (ours) 1609.6±92.6 1475.8±127.2 1353.9±132.3 1409.8±205.8 IBCB (ours) 1263.2±104.3 1262.1±198.3 1287.0±183.3 1257.8±117.3
Algorithm Train Time (sec.) Algorithm Train Time (sec.)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB / / / / SBTS / / / /
B-IRL 41.466±2.898 42.802±10.953 40.348±2.888 42.361±4.737 B-IRL 38.332±4.622 35.703±0.101 35.867±0.287 39.226±4.344
BaseSVC 210.83±37.72 187.31±19.48 181.46±21.53 193.40±14.25 BaseSVC 152.30±13.81 147.73±15.27 146.71±8.414 145.46±7.936
B-ICB-S 90.922±4.373 89.637±4.892 88.908±2.650 90.525±2.571 B-ICB-S 85.615±1.072 85.511±2.658 84.919±1.306 86.802±1.530
IBCB (Ours) 1.370±0.238 1.308±0.198 1.403±0.243 1.191±0.175 IBCB (Ours) 0.889±0.015 0.916±0.028 0.898±0.020 0.937±0.079
TABLE V: Fairness-aware expert policy (SBUCB) parameters’ estimation (Upper Subtable) and Fairness-aware expert policy (SBTS) parameters’ estimation (Lower Subtable) on synthetic dataset. Expert selects action from top-$\bm{2}$ actions with proportion of its learned probability distribution through all actions. Expert policy produces same actions compared with itself, so Batch Test Log Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness Algorithm Batch Test Average Reward
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 1 1 1 1 SBUCB 0.589±0.008 0.583±0.030 0.579±0.029 0.568±0.032
B-IRL 0.445±0.003 0.431±0.007 0.432±0.010 0.430±0.014 B-IRL 0.536±0.012 0.524±0.039 0.519±0.040 0.504±0.045
BaseSVC 0.446±0.004 0.439±0.009 0.439±0.014 0.434±0.015 BaseSVC 0.530±0.013 0.516±0.041 0.511±0.042 0.495±0.047
B-ICB-S 0.431±0.014 0.417±0.008 0.419±0.012 0.421±0.007 B-ICB-S 0.543±0.005 0.512±0.025 0.515±0.024 0.508±0.027
IBCB (Ours) 0.455±0.003 0.461±0.008 0.464±0.008 0.464±0.012 IBCB (Ours) 0.609±0.011 0.600±0.025 0.591±0.030 0.585±0.033
Algorithm Batch Test Log Fitness Algorithm Batch Test Average Reward
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBTS 1 1 1 1 SBTS 0.608±0.023 0.614±0.033 0.612±0.029 0.606±0.027
B-IRL 0.448±0.005 0.444±0.009 0.444±0.006 0.442±0.007 B-IRL 0.558±0.029 0.565±0.044 0.563±0.039 0.555±0.035
BaseSVC 0.451±0.007 0.452±0.007 0.449±0.005 0.449±0.008 BaseSVC 0.553±0.030 0.562±0.047 0.560±0.042 0.552±0.037
B-ICB-S 0.409±0.023 0.396±0.027 0.406±0.017 0.402±0.013 B-ICB-S 0.536±0.022 0.530±0.018 0.541±0.016 0.529±0.022
IBCB (Ours) 0.472±0.006 0.473±0.008 0.473±0.007 0.476±0.003 IBCB (Ours) 0.617±0.024 0.622±0.035 0.621±0.033 0.616±0.028
The experimental results in Table IV and Table V demonstrate that IBCB effectively learns expert parameters in complex, fairness-oriented recommendation settings. With a significantly lower CFR compared to the baseline and well-managed training time, IBCB achieves superior BT-Fitness and BT-AR, showing its capacity to uphold high recommendation quality while ensuring fairness. IBCB’s exceptional performance indicates its robust generalization in learning fair expert parameters and its framework’s adaptability in addressing BCB-related challenges.
5.6 Ablation Study
In this section, we explore extreme scenarios faced by imitation learning (IL) algorithms and IBCB, including delayed updates to training logs, which provides only partial access to the expert’s behavioral history (e.g., the first half of the data), and the emergence of out-of-distribution or contradictory data. Since hyperparameter $\alpha$ of the original expert policy is inaccessible to IBCB, we also examine the impact of different $\alpha$ settings in IBCB. Ablation Study’s results on synthetic dataset are detailed in Section 5.6.1–Section 5.6.5. Also, comprehensive results and analysis of IBCB’s hyperparameter tuning and ablation study on ML-100K dataset are detailed in Section 5.6.6 and Section 5.6.7.
5.6.1 Out-of-distribution Data Experiments on Synthetic Dataset
TABLE VI: Expert policy (SBUCB) parameters’ estimation (Upper Subtable) and Expert policy (SBTS) parameters’ estimation (Lower Subtable) for Out-of-distribution data on synthetic dataset. Expert policy produces same actions compared with itself, so BT-Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness (OOD) Algorithm Batch Test Average Reward (OOD)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBUCB 1 1 1 1 SBUCB 0.709±0.015 0.695±0.036 0.694±0.031 0.684±0.030
B-IRL 0.802±0.005 0.793±0.013 0.795±0.015 0.793±0.016 B-IRL 0.642±0.019 0.622±0.047 0.622±0.042 0.608±0.041
BaseSVC 0.764±0.009 0.756±0.021 0.756±0.022 0.750±0.023 BaseSVC 0.627±0.021 0.607±0.052 0.606±0.046 0.591±0.046
B-ICB-S 0.775±0.039 0.738±0.065 0.755±0.050 0.754±0.025 B-ICB-S 0.632±0.008 0.600±0.030 0.607±0.032 0.596±0.032
IBCB (Ours) 0.970±0.012 0.960±0.036 0.959±0.034 0.969±0.02 IBCB (Ours) 0.717±0.017 0.707±0.034 0.706±0.024 0.693±0.028
Algorithm Batch Test Log Fitness (OOD) Algorithm Batch Test Average Reward (OOD)
std=0 std=0.03 std=0.07 std=0.10 std=0 std=0.03 std=0.07 std=0.10
SBTS 1 1 1 1 SBTS 0.719±0.028 0.739±0.032 0.735±0.032 0.729±0.032
B-IRL 0.792±0.005 0.795±0.010 0.793±0.006 0.793±0.005 B-IRL 0.650±0.034 0.676±0.037 0.671±0.039 0.663±0.037
BaseSVC 0.767±0.015 0.782±0.017 0.780±0.019 0.776±0.014 BaseSVC 0.640±0.041 0.670±0.045 0.666±0.047 0.658±0.044
B-ICB-S 0.714±0.065 0.710±0.096 0.712±0.081 0.724±0.086 B-ICB-S 0.620±0.028 0.643±0.028 0.642±0.026 0.637±0.027
IBCB (Ours) 0.965±0.013 0.962±0.009 0.961±0.003 0.962±0.005 IBCB (Ours) 0.722±0.035 0.744±0.034 0.743±0.033 0.732±0.038
As shown in Table VI, IBCB still outperforms other IL baselines in expert policy’s parameter estimation on out-of-distribution (OOD) data.
5.6.2 Limited Training logs of Experts on Synthetic Dataset
TABLE VII: Reward parameters’ estimation on synthetic dataset (using first halfof expert’s behavior evolution history logs). BaseSVC cannot obtain reward parameters, so it was excluded from OL-Fitness comparison.
Algorithm Online Train Log Fitness (Half Train)
std=0 std=0.03 std=0.07 std=0.10
SBUCB 0.883±0.007 0.879±0.016 0.878±0.018 0.874±0.019
B-IRL 0.752±0.018 0.729±0.051 0.728±0.050 0.721±0.047
B-ICB-S 0.730±0.039 0.727±0.045 0.721±0.063 0.717±0.041
IBCB (Ours) 0.895±0.011 0.894±0.016 0.892±0.024 0.886±0.017
TABLE VIII: Expert policy parameters’ estimation on synthetic dataset (using first half of expert’s behavioral evolution history logs). Expert policy produces same actions compared with itself, so BT-Fitness of expert policy is 1.
Algorithm Batch Test Log Fitness (Half Train)
std=0 std=0.03 std=0.07 std=0.10
@@ 5.6.2 Limited Training logs of Experts on Synthetic Dataset
B-IRL 0.718±0.012 0.699±0.031 0.702±0.035 0.694±0.037
BaseSVC 0.656±0.019 0.638±0.039 0.638±0.043 0.627±0.045
B-ICB-S 0.695±0.054 0.707±0.069 0.695±0.084 0.688±0.043
IBCB (Ours) 0.891±0.019 0.919±0.030 0.890±0.031 0.890±0.023
Algorithm Batch Test Average Reward (Half Train)
std=0 std=0.03 std=0.07 std=0.10